(13%) Problem 4: A stone is dropped from rest from the top of a cliff for it to reach the ground. . ft takes 47 s
> 17% Part (a) what is the initial velocity, vi, of the stone in ms? Grade Summary Deductions 09 Potential 100 Attempts remaining: 5 (0% per attempt) detailed view cotan0asin acos0 atan) acotan)sinh0 coshO tanhO cotanh0 Degrees Radians 0 IND vo BACKSPACE: Delete a token before the cursor position Submit Hint Hints: 0% deduction per hint. Hints remaining:- Feedback: 0% deduction per feedback. 17% Part (b) What is the value of the magnitude of acceleration, in ma?? 17% Part (c) Express the final velocity vr of a ob ect traveling with an acceleration a for a time period At which starts with an initial velocity - 17 % Part (d) Calculate the magnitude of vf in ms. 4m% Part (e) Express the displacement yofano t travelling with acceleration a during a time period A, that has an initial velocity ー▲ 17% Part (n Calculate the magnitude of Ay in meters
Answer #1

Part (a)

Since the stone is dropped and not thrown,initial velocity is zero

ANSWER: ui = Orn/s

======================

Part (b)

There are no external forces acting on the stone, it is falling under influence of gravity

so acceleration is equal to acceleration due to gravity

a=g

ANSWER: a - 9.81m/s a-9.81m/s

==================

Part (c)

Use formula v=u+at

ANSWER: {color{Red} v_{f}=v_{i}+a*Delta t}

===========================

Part (d)

v_{f}=v_{i}+a*Delta t

vfOm/s9.81m/s2 4.7s

ANSWER: 46.107m/s

===========================

Part (e)

Use formula sut +1/2at2

ANSWER: {color{Red} Delta y=v_{i}*Delta t+1/2a*Delta t^{2}}

===========================

Part (f)

Delta y=v_{i}*Delta t+1/2a*Delta t^{2}

.Ay-Orn/s * 4.7s + 1/2 * 9.81m/s2 * (4.7s)

Delta y=0+0.5*9.81*4.7^{2}

ANSWER: {color{Red} Delta y=108.35m}

========================